you read 52 pages per hour
a. you read for an hour in a half.How many pages have you read?
b. You read 130 pages for homework.can you complete the assignment in two hours? Explain your reasoning.

Found 3 solutions by TimothyLamb, Danielshinmath, josmiceli:
Answer by TimothyLamb(4379)

a:
s = q/t
s = 52 pages/hr
q = st
q = 52*1.5
q = 78 pages
---
b:
t = q/s
t = 130/52
t = 2.5 hours
no you cannot, you'll need 2.5 hours to read 130 pages
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by Danielshinmath(31)

#1. You have to do 52/2 because 1 hr/2 = 1/2 hour. So it is 26 pages.
#2. You can only read 52*2=104 pages in 2 hours so you can't complete the assignment.

Answer by josmiceli(19441)

( 52 pages ) / ( 1 hr ) x ( 3/2 hr ) = 78 pages
----------------
In 2 hrs, you are able to read:
( 52 pages ) / ( 1 hr ) x ( 2 hrs ) = 104 pages
130 pages will take more than 2 hrs
Note that +130+-+104+=+26+
26 pages is half of 52, so 130 pages will
take 2 and 1/2 hrs to read